OpenStudy (aurorab):

Solve for x: -3|x - 3| = -6

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

First divide both sides of the equation by -3, so you get /x-3/ = 2. Now, x-3 = 2 or x-3 = -2 If x - 3 = 2, then x = 5; If x - 3 = -2, then x = 1. So you have two solutions for the equation.

OpenStudy (skullpatrol):

@AuroraB Are you still there?

OpenStudy (aurorab):

yeppers!

OpenStudy (skullpatrol):

Did you learn anything?

OpenStudy (skullpatrol):

@AuroraB now you can memorize the answer and hope you get the question on your test.

OpenStudy (skullpatrol):

Have a nice day.